OpenStudy (anonymous):

What are two equivalent fractions for 3/11

OpenStudy (compassionate):

If you multiple the numerator and denominator by the same number they will be equivalent.

OpenStudy (anonymous):

Idk the answer

OpenStudy (anonymous):

3*3/11*3=9/33

OpenStudy (cgreenwade2000):

There are infinite number of answers.

OpenStudy (anonymous):

3*4/11*4=12/44

OpenStudy (cgreenwade2000):

Make her do some work next time jasonxx

OpenStudy (anonymous):

@cgreenwade2000 bro i haven't done anything , don't you remember your answer , you said there are infinite possibilities i've done just two that tends to zero..think again boy and @isabella16 clear?

OpenStudy (cgreenwade2000):

There are infinite number of answers, correct. But by giving her two of them, you've answered the question. And seeing as how she probably doesn't understand how you got them, you provided an answer for her. You could've have just given one example and provided her with something to do. Think again jason.

OpenStudy (anonymous):

ha ha....you haven't stated this earlier, so logically you're incorrect and think what i said by the way i'll keep your point in mind from now on ..thank you